OpenStudy (anonymous):

please help me, can't find the answer anywhere. f^-1(3) if f(x)=3x+ln x I need to find a fractional solution. I think it's 4/5 or something like that, but I'm not sure. If anyone here is adept at Pre-Calculus please respond.

ganeshie8 (ganeshie8):

You want to solve \(x\) in : \(3 = 3x+\ln x\)

OpenStudy (anonymous):

your job is to solve \[3x+\ln(x)=3\]

OpenStudy (anonymous):

don't try algebra do it "by inspection"

OpenStudy (anonymous):

am I supposed to try different numbers in the place of x until I find it?

OpenStudy (anonymous):

i.e guess at a solution make an educated guess you cannot solve this using algebra because you have both \(3x\) and \(\ln(x)\) make a guess i bet you get it on the first or second try

OpenStudy (anonymous):

hint, start with the smallest positive integer

ganeshie8 (ganeshie8):

there is a very natural number to try usually your teacher makes it easy for you to guess, give it a try..

OpenStudy (anonymous):

this is what I have so far. 3=x+e^3

jimthompson5910 (jim_thompson5910):

as @satellite73 said, you'll probably get it on your \(\large \text{first}\) try

OpenStudy (anonymous):

1 works, but he said that's not the answer he's looking for.

OpenStudy (anonymous):

no you are trying to manipulate it (incorrectly) at the risk of repeating myself don't be shy, give it a try

OpenStudy (anonymous):

really?

jimthompson5910 (jim_thompson5910):

x = 1 is the only solution

OpenStudy (anonymous):

yeah, he wants us to solve it and apparently you can find a fractional answer

OpenStudy (anonymous):

cause there aint another one

OpenStudy (anonymous):

he handed out a multiple choice quiz last week with this question and I guessed it, but it's coming up on the test soon and I want to figure it out before. all of the options were fractions.

ganeshie8 (ganeshie8):